# TCAD Simulation of Stitching for Passive CMOS Strip Detectors

### v1

@@ Abstract
This paper shows 3D TCAD simulations of the stitching area to investigate the possible effects stitching can have on the performance of the strip detectors, considering different stitching mismatches. We will show that the mismatch of stitched structures up to does not impact the performance with TCAD simulations which agrees with the results obtained from the measurements.
@@ 1 Introduction
Large area strip detectors are currently being manufactured by very few foundries, limiting the availability, which can be a constraint when building future experiments. CMOS foundries have a long experience building pixelated detectors, starting with the imaging CMOS cameras and exporting the technology for high energy particle experiments. For building large area strip detectors, CMOS foundries have to stitch together several reticles, and the strips have to be stitched between reticles. The impact of stitching reticles during the fabrication is unknown although the implant of the strip might be affected. Pixelated sensors avoid stitching pixel implants but strip detectors have to be stitched in the read out implant.
@@
With the TCAD simulations we investigate the effect of the stitching in the strip sensors, and show that, studying possible mismatching cases up to , stitching strips does not affect perceptively the performance of the strips. These results agree with the measurements shown in previous works [2, 3, 4, 5, 6].
@@ 3.1 Current-Voltage Characteristics
The simulation of the current-voltage characteristic gives an indication of whether a structure leads to an earlier breakdown or to a higher leakage current, which can lead to an increase in the noise of the detector. Figure 4 left plot shows the simulation of the current voltage characteristics for the studied cases up to a bias voltage of . Three of them show the same leakage current as the non-stitched case. while the case with the longitudinal shift stitching shows two order of magnitude higher leakage current than the others. Nevertheless this is a simulation, when scaled to a full reticle (for example, the reticle used for the built LFoundry passive strip detectors) the increase of the leakage current is rather low, as can be appreciated in Figure 4 right plot.

### v2

@@ Abstract
This paper shows 3D TCAD simulations of the stitching area to investigate the possible effects stitching can have on the performance of the strip detectors, considering different stitching mismatches. We will show that the mismatch of stitched structures up to does not impact the performance with TCAD simulations.
@@ 1 Introduction
Large area strip detectors are currently being manufactured by very few foundries, limiting the availability, which can be a constraint when building future experiments. CMOS foundries have a long experience building pixelated detectors, starting with the imaging CMOS cameras and exporting the technology for high energy particle experiments. For building large area strip detectors, CMOS foundries have to stitch together several reticles, and the strip implants and metal layers have to be stitched between reticles. The impact of stitching reticles during the fabrication is unknown although the implant of the strip might be affected. Pixelated sensors avoid stitching pixel implants but strip detectors have to be stitched in the read out implant.
@@
With the TCAD simulations we investigate the effect of the stitching in the strip sensors, and show that, studying possible mismatching cases up to , stitching strips does not affect perceptively the performance of the strips. These results agree with the results shown in previous works [2, 3, 4, 5, 6] where stitching shows no impact on the strip performance.
@@ 3.1 Current-Voltage Characteristics
The simulation of the current-voltage characteristic gives an indication of whether a structure leads to an earlier breakdown or to a higher leakage current, which can lead to an increase in the noise of the detector. Figure 4 left plot shows the simulation of the current voltage characteristics for the studied cases up to a bias voltage of . Three of them show the same leakage current as the non-stitched case. while the case with the longitudinal shift stitching shows two order of magnitude higher leakage current than the others. Nevertheless this is a simulation, when scaled to a full reticle (for example, the reticle used for the built LFoundry passive strip detectors) the increase of the leakage current is rather low, as can be appreciated in Figure 4 right plot. Figure 4 right plot shows the current voltage characteristics simulation for the different mismatches scaled for long reticle.